OpenStudy (taepoohwae):

Find the interger pair that has the given product and sum.The product is 12 The sum is -7. What are the two intergers?

OpenStudy (anonymous):

the necessary concept is that one of the numbers is x and the other number is y. so the product would be. x*y = 12 and the sum is ?

OpenStudy (taepoohwae):

What confuses me is tha it could be several numbers

OpenStudy (anonymous):

integer condition will limit your confuse

OpenStudy (anonymous):

x*y = 12 x+y = -7 // solve the top equation for x or y // plug in x or y into the second equation to find either x or y. then plug x or y back into the first equation to find the second unkown variable.